# Act in Collusion: A Persistent Distributed Multi-Target Backdoor in Federated Learning

### v1

@@ Abstract
Federated learning, a novel paradigm designed to protect data privacy, is vulnerable to backdoor attacks due to its distributed nature. Current research often designs attacks based on a single attacker with a single backdoor, overlooking more realistic and complex threats in federated learning. We propose a more practical threat model for federated learning: the distributed multi-target backdoor. In this model, multiple attackers control different clients, embedding various triggers and targeting different classes, collaboratively implanting backdoors into the global model via central aggregation. Empirical validation shows that existing methods struggle to maintain the effectiveness of multiple backdoors in the global model. Our key insight is that similar backdoor triggers cause parameter conflicts and injecting new backdoors disrupts gradient directions, significantly weakening some backdoors performance. To solve this, we propose a Distributed Multi-Target Backdoor Attack (DMBA), ensuring efficiency and persistence of backdoors from different malicious clients. To avoid parameter conflicts, we design a multi-channel dispersed frequency trigger strategy to maximize trigger differences. To mitigate gradient interference, we introduce backdoor replay in local training to neutralize conflicting gradients. Extensive validation shows that 30 rounds after the attack, Attack Success Rates of three different backdoors from various clients remain above 93%. The code will be made publicly available after the review period.
@@ Appendix G Robustness Validation on Each Dataset
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: See Abstract & Introduction.
  5. 2.
    Limitations
  6. Question: Does the paper discuss the limitations of the work performed by the authors?
  7. Answer: [Yes]
  8. Justification: See Limitations and Ethical Statements.
  9. 3.
    Theory Assumptions and Proofs
  10. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  11. Answer: [Yes]
  12. Justification: See Methodology & Appendix.
  13. 4.
    Experimental Result Reproducibility
  14. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  15. Answer: [Yes]
  16. Justification: See Abstract & Section 4.1.
  17. 5.
    Open access to data and code
  18. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  19. Answer: [Yes]
  20. Justification: See Abstract & Section 4.1. Our core code is uploaded in a zip package.
  21. 6.
    Experimental Setting/Details
  22. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  23. Answer: [Yes]
  24. Justification: See Section 4.1.
  25. 7.
    Experiment Statistical Significance
  26. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  27. Answer: [Yes]
  28. Justification: See Section 4.2.
  29. 8.
    Experiments Compute Resources
  30. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  31. Answer: [Yes]
  32. Justification: See Section 4.1.
  33. 9.
    Code Of Ethics
  34. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  35. Answer: [Yes]
  36. Justification: All aspects of our research adhere to the NeurIPS Code of Ethics.
  37. 10.
    Broader Impacts
  38. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  39. Answer: [Yes]
  40. Justification: See Ethical Statements.
  41. 11.
    Safeguards
  42. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  43. Answer: [Yes]
  44. Justification: See Ethical Statements.
  45. 12.
    Licenses for existing assets
  46. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  47. Answer: [Yes]
  48. Justification: We clearly mention the origin of all assets used, respect intellectual property rights and comply with the terms of the relevant licences.
  49. 13.
    New Assets
  50. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  51. Answer: [Yes]
  52. Justification: We have specifically added the README file to the submitted code.
  53. 14.
    Crowdsourcing and Research with Human Subjects
  54. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  55. Answer: [N/A]
  56. Our study does not address that question.
  57. 15.
    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
  58. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  59. Answer: [N/A]
  60. Our study does not address that question.

### v2

@@ Abstract
Federated learning, a novel paradigm designed to protect data privacy, is vulnerable to backdoor attacks due to its distributed nature. Current research often designs attacks based on a single attacker with a single backdoor, overlooking more realistic and complex threats in federated learning. We propose a more practical threat model for federated learning: the distributed multi-target backdoor. In this model, multiple attackers control different clients, embedding various triggers and targeting different classes, collaboratively implanting backdoors into the global model via central aggregation. Empirical validation shows that existing methods struggle to maintain the effectiveness of multiple backdoors in the global model. Our key insight is that similar backdoor triggers cause parameter conflicts and injecting new backdoors disrupts gradient directions, significantly weakening some backdoors performance. To solve this, we propose a Distributed Multi-Target Backdoor Attack (DMBA), ensuring efficiency and persistence of backdoors from different malicious clients. To avoid parameter conflicts, we design a multi-channel dispersed frequency trigger strategy to maximize trigger differences. To mitigate gradient interference, we introduce backdoor replay in local training to neutralize conflicting gradients. Extensive validation shows that 30 rounds after the attack, Attack Success Rates of three different backdoors from various clients remain above 93%.